PUMP-PROBE DELAYED IONIZATION STUDY OF PHENYLACETYLENE

\underline{GARY V. LOPEZ}, PHILIP M. JOHNSON AND TREVOR J. SEARS, Department of Chemistry, Stony Brook University, Stony Brook, New York 11794; CHIH-HSUAN CHANG, Department of Chemistry, Brookhaven National Laboratory, Upton, New York 11973.

Pump-probe photoionization experiments on the origin of the S_{1} ← S_{0} transition of phenylacetylene (PA) have revealed the production of long lived (≫ 100 μsec lifetime) species with low ionization potential where short lifetimes (nsec lifetime) are expected if simple intersystem crossing takes place.^{a} The pump-probe delay photoionization decay curve for the band origin, previously presented by Hofstein et al.,^{a} and four other assigned strong bands in the S_{1} ← S_{0} transition of PA have been obtained in recent experiments. Following S_{1} ← S_{0} excitation, the decay consists of a prompt short-lived component matching the measured singlet fluorescence lifetime and a second, very long-lived, component. The ratios between the singlet channel signal to that of the long-lived species for these bands are not the same: 3.03±0.38, 1.48±0.33, 1.75±0.19, 3.11±0.36 and 2.77±0.35 for the origin, ν_{15}, ν_{6a}, β_{cc} and ν_{1}, respectively. The fact that the ν_{15} and ν_{6a} modes have low ratios suggests that the normal coordinates of these modes could lie near the reaction path for the formation of the long lived species. Thus, these results combined with theoretical calculations may potentially give an indication of the identity of the long-lived species. Acknowledgments: Work at Brookhaven National Laboratory was carried out under Contract No. DE-AC02-98CH10886 with the U.S. Department of Energy and supported by its Office of Basic Energy Sciences, Division of Chemical Sciences, Geosciences and Biosciences.

aJ. Hofstein, H. Xu, T. J. Sears and P. M. Johnson, J. Phys. Chem. A 112, 1195-1201 (1999)